OpenStudy (anonymous):

SOLVE FOR X: 3/5 s = 60

OpenStudy (anonymous):

when u write this problem there is a difference between (3/5)s=60 and 3/(5s)=60 so u have to b clear

OpenStudy (dumbcow):

Multiply by reciprocal on both sides \[s = 60*\frac{5}{3}\]

OpenStudy (anonymous):

3/5 or 5/3??